# The Cosmological CPT Theorem

### v2

@@ 7.2 The Loop Momentum Integral
[…]
@@
[…]
@@ 7.3 Spinning Fields
[…]
@@
[…]
@@
[…]
@@ 7.4 Constraints on Boundary Wavefunction Coefficients
where for the case of scalars, the mass $m^{2}/H^{2}=m^{2}\ell^{2}=\Delta(d-\Delta)$ is related to the conformal dimension in the usual way $\Delta=d/2+\nu$, $\nu=\sqrt{d^{2}/4-m^{2}/H^{2}}\,$. [^59] Note for AdS this relation is $m^{2}\ell_{AdS}^{2}=\Delta(\Delta-d)$. In AdS $\Delta^{\pm}\in\mathbb{R}$ for scalar fields of any mass. In dS for scalar fields with mass $m^{2}/H^{2}>d^{2}/4$, $\Delta^{\pm}\in\mathbb{C}$ which is known as the principal series; fields with mass $m^{2}/H^{2}\leq d^{2}/4$, $\Delta^{\pm}\in\mathbb{R}$ which is known as the complementary series. For massive $s$ fields the mass and spin are related to the conformal dimension by $\Delta=d/2+\mu$, $\mu=\sqrt{(d+2s-4)^{2}/4-m^{2}/H^{2}}\,$. We define $\overline{\phi}_{+}$ with $\Delta^{+}\equiv\Delta$ and $\overline{\phi}_{-}$ (with $\Delta^{-}=d-\Delta$). For heavy scalar fields ($m>dH/2$), $\Delta$ is complex and so these boundary operators do not represent a self-adjoint basis [^60] Of course, we could have made the choice to expand the fields in a self-adjoint way but this would sacrifice boundary conformal invariance as such an expansion would necessarily mix terms of different weights. See e.g. [76, 77, 78, 79, 80, 81, 82, 83, 84, 85] for discussions regarding the principal series.. We will therefore restrict our discussion to light fields and leave a more comprehensive examination, including heavy fields, to future work. In this case, $\Delta$ is real and positive so the $\overline{\phi}_{-}$ term dominates. Starting from the expansion of the wavefunction in terms of its coefficients, (3.73), we can similarly define a “boundary” wavefunction [^61] Although this is referred to as a “boundary” wavefunction in the literature, this is not really a wavefunction living at the $\mathcal{I}^{+}$ where $\eta=0^{-}$, since from (7.77) it is clear that only massless fields would survive in the limit of $\eta=0^{-}$. The more accurate description is that the coefficients of the wavefunction have had their $\eta$-dependence stripped off. which is a functional of these boundary fields $\overline{\phi}_{-}$ and has no explicit $\eta$-dependence
@@
Let us now count the scalings of each term in the exponent one by one to see how $\overline{\psi}_{n}(\textbf{k}_{a})$ scales under $\to\lambda\textbf{k}$ k which in turn will scale the internal and external energies $k\to\lambda k$. The $d$-dimensional $\mathop{}\!\mathrm{d}^{d}k$ measures each scale with volume and thus they will scale with an overall factor of $\lambda^{nd}$. This cancels with an equivalent scaling of the Fourier transformed $\overline{\phi}_{-}(\textbf{k}_{a})$ which give an overall scaling of $\lambda^{-nd}$. [^64] This is in the exact same way that in (3.73) the original bulk field $\phi_{\textbf{k}_{a}}$ scaled as $\lambda^{d}$ resulting in the bulk wavefunction coefficient scaling inversely to the $\delta^{d}\left(\sum\textbf{k}_{a}\right)$. Each boundary field must contribute an extra scaling like $\eta^{\Delta-d}$ to ensure the Scale Invariance of the bulk field (7.78). The result of this is that in addition to the factor coming from the Fourier Transform the fields contribute a further $\lambda^{dn-\sum_{\alpha}\Delta_{\alpha}}$. Finally, the $\delta^{d}(\textbf{k}_{1}+\ldots+\textbf{k}_{n})$ scales with inverse volume $\lambda^{-d}$. [^65] Usually in the context of computations in de Sitter, the $\delta({\textbf{k}_{1}}+\ldots+{\textbf{k}_{n}})$-function is omitted but we remind the reader of its importance in terms of arguments involving Scale Invariance, as it scales with inverse volume, as well as its necessity for ensuring momentum conservation which comes from translation invariance in position space. We know that the exponent of (7.79) must be dimensionless, thus by dimensional analysis, any IR-finite boundary wavefunction coefficient will scale as
@@
$\bf{RR}$ does not care about the time at which the wavefunction is evaluated and so the resulting constraint is unchanged: [^66] We remind the reader that the conventions we use in cosmology where the conformal time on conformally flat Poincaré slicing $\eta\in[-\infty,0^{-}]$ is negative in the expanding branch and increases from $\infty$- ∞ in the infinite past to $\mathcal{I}^{+}=0^{-}$.
[…]
@@
[…]
@@
[…]
@@
The steps leading up to this phase formula implicitly assumed that the amplitudes were IR and UV-finite. We have already discussed the $-i\pi$ shift due to UV $\log$ divergences in Section 6.2; we now make the corresponding remark for IR divergences. For a typical IR-divergent amplitude, there is a term like $C\log(-\eta)$ which, upon rotating $\eta$ from $0^{-}$ to $0^{+}$, becomes $C\log(-e^{-i\pi}\eta)$ This provides an extra $-i\pi C$ shift that adjusts the IR-finite piece of the amplitude, which does not conform to (7.94). But, the leading order IR-divergence will continue to satisfy (7.94). An example of such a calculation will be quoted in Section 7.5 below.
@@ 7.5 Explicit Checks of the Phase Formula
The first five calculations iv below all involve conformally coupled scalars (denoted by $\varphi$), while the remaining two vi and vii involve massless fields (denoted by $\phi$). For the calculations in iiivii, we work in non-integer dimensions for the purposes of dimensional regularisation (dim-reg) using the prescription described in Appendix C of [28]. In this approach, the mass of the field is also renormalised to keep the order of the Hankel function $\nu$ fixed, which ensures that the integrals can be computed analytically in the dim-reg parameter $\delta$ [^69] An alternative prescription proposed in [134] does not renormalise the mass of the field, resulting in the order of the Hankel function for the case of massless fields in $d=3+\delta$-dimensions becoming $\nu=d/2=(3+\delta)/2$. However, this prevents the integrals from being computed analytically. Consequently, the authors carry out an expansion in $\delta$ prior to computing the integrals which introduces non-trivial corrections at various points in the calculation. This makes the overall answer non-analytic in $\delta$ and prevents a direct comparison with the phase formula in (7.93) and (7.94).
@@
The phase of the contact three-point wavefunction coefficient $\overline{\psi}^{\text{tree}}_{3,\varphi\varphi\tilde{\sigma}}$, generated via a simple cubic interaction $\varphi\varphi\tilde{\sigma}$ in 3+13+13 + 1-dimensions, involving two conformally-coupled scalars (with $\Delta=2$) and a massive scalar $\tilde{\sigma}$ with mass $m<\sqrt{2}H$ (corresponding to $2<\Delta\leq 3$), in order for the time integral to converge in the IR (i.e. when $\eta=0^{-}$), is found in Appendix B of [13] to be
@@
The contact three-point wavefunction coefficient $\overline{\psi}^{\text{tree}}_{3,\varphi\varphi\varphi}$ for the interaction involving three conformally-coupled scalars, generated via a simple cubic interaction $\varphi\varphi\varphi$ in 3+13+13 + 1-dimensions, where the time integral diverges in the IR (i.e. when $\eta=0^{-}$), is found to be
@@
iv. Conformal 111-loop UV-finite coefficient in non-integer dimension $d=3+\delta$,
@@
The phase of the 111-site 111-loop four-point wavefunction coefficient $\overline{\psi}^{(L=1)}_{4,\varphi^{6}}$ for the interaction involving four conformally-coupled scalars, generated via a 6th-order interaction $\varphi^{6}$ in non-integer $d=3+\delta$-dimensions (which, as previously explained, rescales the conformal dimension of the conformally coupled scalar $\Delta=d/2+1/2=(4+\delta)/2$), was computed in [135]. However, an incorrect phase was found, which will be corrected in [93]. For this particular Feynman-Witten diagram one finds
@@
v. Conformal 111-loop UV-divergent coefficient in non-integer dimension $d=3+\delta$,
@@
The 222-site, 111-loop de Sitter wavefunction coefficient for arbitrary number $n$ of conformally coupled fields in external legs in non-integer $d=3+\delta$-dimensions will be computed in upcoming work [136, 137] [^70] We thank Tom Westerdijk for kindly sharing these results and related with us [136, 137, 138]. See also similar results in [139]. where the author generalises the recently developed differential equation approach [140, 141] to computing flat FLRW cosmological wavefunction coefficients to loop diagrams. The result for the phase of the UV-divergent term is
@@
vi. Massless 111-loop UV-divergent coefficient in non-integer dimension $d=3+\delta$,
@@
As previously discussed, in cases with a logarithmic UV-divergence, one can renormalise these using dim-reg. In dim-reg, the logarithmic UV-divergence turns into a $1/\delta$ divergence (for IR-finite 111-loop diagrams), which can be canceled by expanding the phase using the simple fact that any generic complex number $A$ can be expressed as $A=|A|e^{i\arg(A)}$. For IR-finite 111-loop diagrams we thus find
@@
This phenomenon was found in [135], where the authors computed a parity-odd contribution to the scalar trispectrum using the in-in formalism. We have demonstrated that this is in fact a generic feature of UV-divergent loop diagrams and will be explored further in [88].
In [28], the 111-loop four-point wavefunction coefficient $\overline{\psi}^{(L=1)}_{2,\phi^{\prime 3}}$ for the interaction involving three massless scalars, generated via a boost-breaking derivative interaction $\phi^{\prime 3}$ in non-integer $d=3+\delta$-dimensions (which rescales the conformal dimension of the massless scalar $\Delta=d/2+3/2=(6+\delta)/2$), was computed with an incorrect phase that will be corrected in [93]. For this particular Feynman-Witten diagram one finds
@@
[…]
@@
Hence for the case of $d=2$ we expect the central charge, which is equivalent to the coefficient of the 2-point stress tensor correlator, to be imaginary at tree level and real at 111-loop order and so on. This matches what has been found in the literature — refer to Section 1.3 for a list of previous results regarding the central charge in dS/CFT.
In fact, the phase formula offers a natural explanation for how a complex de Sitter wavefunction $\Psi$ — with Lorentzian bulk evolution — can be reconstructed from its modulus-squared $|\Psi|^{2}\equiv\Psi\Psi^{*}$, which is typically associated with a doubled Euclidean CFT lacking phase information. For example, in $d=2$ pure gravity, this leads to a theory with $c=26$. Supposing one is given only the probability distribution $|\Psi|^{2}$, the phase formula provides precisely the information needed to define a consistent “square root” $\Psi\equiv Z_{\text{CFT}}$, restoring the correct complex structure of the wavefunction. This suggests that, in principle, nearly any $c=26$ two-dimensional CFT can be used to reconstruct a candidate dS/CFT wavefunction. This perspective also lends support to the recently proposed duality between $dS_{3}$ and the complex Liouville string theory with $c=i\frac{3\ell}{2G_{N}}+13+\mathcal{O}(G_{N})$ [142, 143, 144, 145, 146, 147].

### v3

@@ 11 $\bf{CRT}$ $\bf{RR}$
$\text{$\bf{CRT}$}\left(\psi_{n}^{(L)}\right)$ $=\left[\psi^{(L)}_{n}(e^{-i\pi}\eta_{0},\textbf{k},e^{i\pi}\varepsilon)\right]^{*}=e^{i\pi(d+1)(L-1)}\psi_{n}(\eta_{0};\textbf{k};\varepsilon)\,$, (7.63)
@@
$\text{$\bf{RR}$}\left(\psi_{n}^{(L)}\right)$ $=\left[\psi^{(L)}_{n}(\eta_{0};e^{-i\pi}\textbf{k};\varepsilon)\right]^{*}=e^{i\pi\left((d+1)L-1\right)}\psi_{n}(\eta_{0};\textbf{k};\varepsilon)\,$, (7.65)
@@ 7.3 Spinning Fields
The totally-symmetric, traceless tensor $\sigma_{i_{1}\dots i_{s}}$ has spatial indices $i_{1}=1,\ldots,d$ which span the $d$-dimensional spacelike hypersurface orthogonal to the $\eta$ coordinate. [^57] The Effective Field Theory of Inflation (EFToI) [149] is derived by considering a theory which is only invariant under spatial diffeomorphisms, for which there is a preference for the co-ordinate choice used in unitary gauge, where the time coordinate is chosen to coincide with the surfaces of constant value of the field $\sigma_{i_{1}\dots i_{s}}$. The EFToI thus encapsulates generic class of models of inflation where spatial diffeomorphisms are preserved. (7.69) can be written in a covariant way by using the Goldstone boson $\pi$ of time translations to upgrade the spatial tensor $\sigma_{i_{1}\dots i_{s}}$ to a covariant spacetime tensor. The coupling of $\sigma_{i_{1}\dots i_{s}}$ to $\pi$ is also dictated by this constructions but we will not need this here. $\sigma_{i_{1}\dots i_{s}}$ has $(2s+1)$ components, which each create states (“particles”) with helicities $0,\pm 1,\dots,\pm s$, and we have enforced invariance under dilatations by including inverse factors of the scale factor for each coordinate derivative. Following [15, 148] we Fourier transform and diagonalise this using the helicity modes, $\sigma_{h}$, defined by:
@@
$\textbf{e}^{h}_{i}(\textbf{k})\left[\textbf{e}^{h^{\prime}}_{i}(\textbf{k})\right]^{*}-4\delta_{hh^{\prime}}$ $=0$ $\text{(orthogonality and normalisation)}\,$, (7.72)
$\left[\textbf{e}_{i}^{h}(\textbf{k})\right]^{*}-\textbf{e}_{i}^{h}(-\textbf{k})$ $=0$ $\text{($\sigma_{i_{1}\ldots i_{s}}(x)$ is real)}\,.$ (7.73)
@@
$\psi^{h_{1}\ldots h_{n}}_{n}=\text{(tensor structure)}\times\text{(trimmed wavefunction coefficient)}\,.$ (7.75)
@@ 7.4 Constraints on Boundary Wavefunction Coefficients
where for the case of scalars, the mass $m^{2}/H^{2}=m^{2}\ell^{2}=\Delta(d-\Delta)$ is related to the conformal dimension in the usual way $\Delta=d/2+\nu$, $\nu=\sqrt{d^{2}/4-m^{2}/H^{2}}\,$. [^59] Note for AdS this relation is $m^{2}\ell_{AdS}^{2}=\Delta(\Delta-d)$. In AdS $\Delta^{\pm}\in\mathbb{R}$ for scalar fields of any mass. In dS for scalar fields with mass $m^{2}/H^{2}>d^{2}/4$, $\Delta^{\pm}\in\mathbb{C}$ which is known as the principal series; fields with mass $m^{2}/H^{2}\leq d^{2}/4$, $\Delta^{\pm}\in\mathbb{R}$ which is known as the complementary series. For massive spin-$s$ fields the mass and spin are related to the conformal dimension by $\Delta=d/2+\mu$, $\mu=\sqrt{(d+2s-4)^{2}/4-m^{2}/H^{2}}\,$. We define $\overline{\phi}_{+}$ with $\Delta^{+}\equiv\Delta$ and $\overline{\phi}_{-}$ (with $\Delta^{-}=d-\Delta$). For heavy scalar fields ($m>dH/2$), $\Delta$ is complex and so these boundary operators do not represent a self-adjoint basis [^60] Of course, we could have made the choice to expand the fields in a self-adjoint way but this would sacrifice boundary conformal invariance as such an expansion would necessarily mix terms of different weights. See e.g. [76, 77, 78, 79, 80, 81, 82, 83, 84, 85] for discussions regarding the principal series.. We will therefore restrict our discussion to light fields and leave a more comprehensive examination, including heavy fields, to future work. In this case, $\Delta$ is real and positive so the $\overline{\phi}_{-}$ term dominates. Starting from the expansion of the wavefunction in terms of its coefficients, (3.73), we can similarly define a “boundary” wavefunction [^61] Although this is referred to as a “boundary” wavefunction in the literature, this is not really a wavefunction living at the $\mathcal{I}^{+}$ where $\eta=0^{-}$, since from (7.77) it is clear that only massless fields would survive in the limit of $\eta=0^{-}$. The more accurate description is that the coefficients of the wavefunction have had their $\eta$-dependence stripped off. which is a functional of these boundary fields $\overline{\phi}_{-}$ and has no explicit $\eta$-dependence
@@
Let us now count the scalings of each term in the exponent one by one to see how $\overline{\psi}_{n}(\textbf{k}_{a})$ scales under $\textbf{k}\to\lambda\textbf{k}$ which in turn will scale the internal and external energies $k\to\lambda k$. The $d$-dimensional $\mathop{}\!\mathrm{d}^{d}k$ measures each scale with volume and thus they will scale with an overall factor of $\lambda^{nd}$. This cancels with an equivalent scaling of the Fourier transformed $\overline{\phi}_{-}(\textbf{k}_{a})$ which give an overall scaling of $\lambda^{-nd}$. [^64] This is in the exact same way that in (3.73) the original bulk field $\phi_{\textbf{k}_{a}}$ scaled as $\lambda^{d}$ resulting in the bulk wavefunction coefficient scaling inversely to the $\delta^{d}\left(\sum\textbf{k}_{a}\right)$. Each boundary field must contribute an extra scaling like $\eta^{\Delta-d}$ to ensure the Scale Invariance of the bulk field (7.78). The result of this is that in addition to the factor coming from the Fourier Transform the fields contribute a further $\lambda^{dn-\sum_{\alpha}\Delta_{\alpha}}$. Finally, the $\delta^{d}(\textbf{k}_{1}+\ldots+\textbf{k}_{n})$ scales with inverse volume $\lambda^{-d}$. [^65] Usually in the context of computations in de Sitter, the $\delta({\textbf{k}_{1}}+\ldots+{\textbf{k}_{n}})$-function is omitted but we remind the reader of its importance in terms of arguments involving Scale Invariance, as it scales with inverse volume, as well as its necessity for ensuring momentum conservation which comes from translation invariance in position space. We know that the exponent of (7.79) must be dimensionless, thus by dimensional analysis, any IR-finite boundary wavefunction coefficient will scale as
@@
$\bf{RR}$ does not care about the time at which the wavefunction is evaluated and so the resulting constraint is unchanged: [^66] We remind the reader that the conventions we use in cosmology where the conformal time on conformally flat Poincaré slicing $\eta\in[-\infty,0^{-}]$ is negative in the expanding branch and increases from -$\infty$ in the infinite past to $\mathcal{I}^{+}=0^{-}$.
$\text{$\bf{RR}$}:\left[\overline{\psi}^{(L)}_{n}(\textbf{k})\right]^{*}$ $=e^{i\pi[(d+1)L-1]}\overline{\psi}^{(L)}_{n}(e^{-i\pi}\textbf{k})\,.$ (7.82)
@@
$\textbf{CRT}:$ $\overline{\psi}^{(L)}_{n}(\textbf{k})\!$ $=\,\,e^{-i\pi[(d+1)(L-1)+dn-\sum_{\alpha}\Delta_{\alpha}]}\,\left[\overline{\psi}^{(L)}_{n}(\textbf{k})\right]^{*}\,$, (7.90)
@@
$\textbf{RR}:$ $\,\overline{\psi}^{(L)}_{n}(\textbf{k})\!$ $=\,\,e^{-i\pi[(d+1)L-1]}\,\left[\overline{\psi}^{(L)}_{n}(e^{-i\pi}\textbf{k})\right]^{*}\,$, (7.92)
@@
The steps leading up to this phase formula implicitly assumed that the amplitudes were IR and UV-finite. We have already discussed the $i\pi$ shift due to UV $\log$ divergences in Section 6.2; we now make the corresponding remark for IR divergences. For a typical IR-divergent amplitude, there is a term like $C\log(-\eta)$ which, upon rotating $\eta$ from $0^{-}$ to $0^{+}$, becomes $C\log(-e^{-i\pi}\eta)$ This provides an extra $i\pi C$ shift that adjusts the IR-finite piece of the amplitude, which does not conform to (7.94). But, the leading order IR-divergence will continue to satisfy (7.94). An example of such a calculation will be quoted in Section 7.5 below.
@@ 7.5 Explicit Checks of the Phase Formula
The first five calculations iv below all involve conformally coupled scalars (denoted by $\varphi$), while the remaining two vi and vii involve massless fields (denoted by $\phi$). For the calculations in iiivii, we work in non-integer dimensions for the purposes of dimensional regularisation (dim-reg) using the prescription described in Appendix C of [28]. In this approach, the mass of the field is also renormalised to keep the order of the Hankel function $\nu$ fixed, which ensures that the integrals can be computed analytically in the dim-reg parameter $\delta$. [^69] An alternative prescription proposed in [150] does not renormalise the mass of the field, resulting in the order of the Hankel function for the case of massless fields in $d=3+\delta$-dimensions becoming $\nu=d/2=(3+\delta)/2$. However, this prevents the integrals from being computed analytically. Consequently, the authors carry out an expansion in $\delta$ prior to computing the integrals which introduces non-trivial corrections at various points in the calculation. This makes the overall answer non-analytic in $\delta$ and prevents a direct comparison with the phase formula in (7.93) and (7.94).
@@
The phase of the contact three-point wavefunction coefficient $\overline{\psi}^{\text{tree}}_{3,\varphi\varphi\tilde{\sigma}}$, generated via a simple cubic interaction $\varphi\varphi\tilde{\sigma}$ in $3+1$-dimensions, involving two conformally-coupled scalars (with $\Delta=2$) and a massive scalar $\tilde{\sigma}$ with mass $m<\sqrt{2}H$ (corresponding to $2<\Delta\leq 3$), in order for the time integral to converge in the IR (i.e. when $\eta=0^{-}$), is found in Appendix B of [13] to be
@@
The contact three-point wavefunction coefficient $\overline{\psi}^{\text{tree}}_{3,\varphi\varphi\varphi}$ for the interaction involving three conformally-coupled scalars, generated via a simple cubic interaction $\varphi\varphi\varphi$ in $3+1$-dimensions, where the time integral diverges in the IR (i.e. when $\eta=0^{-}$), is found to be
@@
iv. Conformal $1$-loop UV-finite coefficient in non-integer dimension $d=3+\delta$,
@@
The phase of the $1$-site 11-loop four-point wavefunction coefficient $\overline{\psi}^{(L=1)}_{4,\varphi^{6}}$ for the interaction involving four conformally-coupled scalars, generated via a 6th-order interaction $\varphi^{6}$ in non-integer $d=3+\delta$-dimensions (which, as previously explained, rescales the conformal dimension of the conformally coupled scalar $\Delta=d/2+1/2=(4+\delta)/2$), was computed in [151]. However, an incorrect phase was found, which will be corrected in [93]. For this particular Feynman-Witten diagram one finds
@@
v. Conformal $1$-loop UV-divergent coefficient in non-integer dimension $d=3+\delta$,
@@
The $2$-site, 11-loop de Sitter wavefunction coefficient for arbitrary number $n$ of conformally coupled fields in external legs in non-integer $d=3+\delta$-dimensions will be computed in upcoming work [152, 153] [^70] We thank Tom Westerdijk for kindly sharing these results and related with us [152, 153, 154]. See also similar results in [155]. where the author generalises the recently developed differential equation approach [156, 157] to computing flat FLRW cosmological wavefunction coefficients to loop diagrams. The result for the phase of the UV-divergent term is
@@
vi. Massless $1$-loop UV-divergent coefficient in non-integer dimension $d=3+\delta$,
@@
As previously discussed, in cases with a logarithmic UV-divergence, one can renormalise these using dim-reg. In dim-reg, the logarithmic UV-divergence turns into a $1/\delta$ divergence (for IR-finite $1$-loop diagrams), which can be canceled by expanding the phase using the simple fact that any generic complex number $A$ can be expressed as $A=|A|e^{i\arg(A)}$. For IR-finite $1$-loop diagrams we thus find
@@
This phenomenon was found in [151], where the authors computed a parity-odd contribution to the scalar trispectrum using the in-in formalism. We have demonstrated that this is in fact a generic feature of UV-divergent loop diagrams and has been explored further in [88].
In [28], the $1$-loop four-point wavefunction coefficient $\overline{\psi}^{(L=1)}_{2,\phi^{\prime 3}}$ for the interaction involving three massless scalars, generated via a boost-breaking derivative interaction $\phi^{\prime 3}$ in non-integer $d=3+\delta$-dimensions (which rescales the conformal dimension of the massless scalar $\Delta=d/2+3/2=(6+\delta)/2$), was computed with an incorrect phase that will be corrected in [93]. For this particular Feynman-Witten diagram one finds
@@
where $S_{d-2}$ is the surface area of the (d2)(d-2)-dimensional unit sphere (i.e. $S_{1}=2\pi$) and $I(\delta)$ is a dimensionless quantity with no contribution to the phase defined in  [28]. As we stated in Sections 7.2 and 7.4 the phase formula still holds for the case of derivative interactions; the phase is also important for the consistency of the cutting rules derived in [28].
@@
Hence for the case of $d=2$ we expect the central charge, which is equivalent to the coefficient of the 2-point stress tensor correlator, to be imaginary at tree level and real at $1$-loop order and so on. This matches what has been found in the literature — refer to Section 1.3 for a list of previous results regarding the central charge in dS/CFT.
In fact, the phase formula offers a natural explanation for how a complex de Sitter wavefunction $\Psi$ — with Lorentzian bulk evolution — can be reconstructed from its modulus-squared $|\Psi|^{2}\equiv\Psi\Psi^{*}$, which is typically associated with a doubled Euclidean CFT lacking phase information. For example, in $d=2$ pure gravity, this leads to a theory with $c=26$. Supposing one is given only the probability distribution $|\Psi|^{2}$, the phase formula provides precisely the information needed to define a consistent “square root” $\Psi\equiv Z_{\text{CFT}}$, restoring the correct complex structure of the wavefunction. [^71] Although, not every $c=26$ two-dimensional CFT can be used to reconstruct a candidate dS/CFT wavefunction with a semiclassical gravity dual, as the factorisation into two CFTs, each with a large imaginary part in $c$, is necessary to have a weakly coupled graviton in the bulk dual. This happens because the $c=26$ dS/CFT has two different stress-tensors which mix with each other. This perspective also lends support to the recently proposed duality between $dS_{3}$ and the complex Liouville string theory with $c=i\frac{3\ell}{2G_{N}}+13+\mathcal{O}(G_{N})$ [158, 159, 160, 161, 162, 163].